OpenStudy (anonymous):

1.Solve for the variable. y + -4 = -15 A. -19 B. -11 C. 19 D. 11 2.Which ratio is not equal to 49/21? A. 7/10 B. 7/3 C. 98/42 D. 245/105

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

1. is B but 2 i need help in

jimthompson5910 (jim_thompson5910):

check each possible answer

jimthompson5910 (jim_thompson5910):

49/21 = 7/10 ... is this true?

jimthompson5910 (jim_thompson5910):

49/21 = 7/3 ... is this true? etc etc

OpenStudy (anonymous):

that guy gave me the answer :(

jimthompson5910 (jim_thompson5910):

well hopefully you know how to check for the answer

OpenStudy (anonymous):

divide that bottom number with 21 and see if it goes into it even :)

jimthompson5910 (jim_thompson5910):

well that's one way, but what I had in mind was this 49/21 = 7/3 49*3 = 21*7 ... cross multiply 147 = 147 ... multiply since the last equation is true, the first equation is true